OpenStudy (anonymous):

How can I add 4/9 + 5/6,using LCM? I need to see all the steps. I'm still getting confused once I find the LCM. And I am also trying to multiply 3/4 X 1/8. Can someone help me and explain the steps?

OpenStudy (swissgirl):

uh oh lol

OpenStudy (anonymous):

Break 9 up into its factors to get LCM. 9 is 3 x 3. Do the same with 6. 3 x 2. The LCM has to have everything in both. So, LCM is 3 x 3 x 2. Do you follow so far?

OpenStudy (swissgirl):

we only did it with GCM btw never with LCM

OpenStudy (anonymous):

oop.. it doesn't say to do that.. I just thought I was suppose to do that every time . The problem is just to add 4/9 + 5/6

OpenStudy (swissgirl):

Nooo wait we did it with LCM. Sorry I am confusing u

OpenStudy (swissgirl):

We need to list the multiples of 9 and 6 9=9,18,27,36... U list the multiples of 6 ok?

OpenStudy (anonymous):

Now that you have the LCM = 18, you have to see what you will have had to do to each fraction to get the LCM in the denominator. With 4/9, you have to multiply the numerator and the denominator by 2. That is you multiply by 2/2 which is the same as 1/1 or 1 and you will still have 4/9 in essence because 1 is the multiplicative identity. So, 4/9 x 2/2 = 8/18. Still with me?

OpenStudy (anonymous):

I think I'm the one whose confused here. sorry.. the problem is just for me to add the fractions....

OpenStudy (swissgirl):

U can only add the fractions if the have the same denomanator and in this case they dont one is 9 and the other is 6 we need to have the same denomanator

OpenStudy (anonymous):

Yes, and you do have get the LCM to add those fractions, so we continue. 5/6 = 5/6 x 3/3 = 15/18. Notice how both fractions now have the same denominator, 18.

OpenStudy (anonymous):

9=18,27,36,45,54.... 6=12,24,30,36,,42,

OpenStudy (anonymous):

Now, 8/18 + 15/18 = 23/18 since with fractions you keep the denominator the same and just add the numerators.

OpenStudy (swissgirl):

6=6,12,18,24,30,36

OpenStudy (swissgirl):

u skipped out the 18

OpenStudy (anonymous):

So, your answer is 23/18 and you are DONE unless you want to simplify that answer. That's all there is to it.

OpenStudy (anonymous):

Do you understand the steps I've given?

OpenStudy (anonymous):

no, you went to fast and I got lost

OpenStudy (swissgirl):

ok well we have the multiples of both what is the Lowest common multiple?

OpenStudy (anonymous):

Just go over it, it's all there or if you like, ask a question.

OpenStudy (anonymous):

Its too hard for me to try to follow both of you. sorry

OpenStudy (anonymous):

To do LCM, you can list the multiples of each denominator and take the smallest of the common numbers. That is one way to do it and is a little slow. The faster way is to list the factors of each and make sure that the LCM has all factors from both.

OpenStudy (anonymous):

Swissgirl, what do u mean I skipped out 18

OpenStudy (swissgirl):

6=12,24,30,36,,42, When u listed out the multiples of 6 u skipped out 18 It should have been 6=12,18,24,30,36,,42,

OpenStudy (anonymous):

sorry

OpenStudy (swissgirl):

and u skipped out 6 cuz 6*1=6

OpenStudy (anonymous):

ok

OpenStudy (swissgirl):

So the multiples of 6 are 6=6,12,18,24,30,36,42...

OpenStudy (anonymous):

9=18,27,36,45,54 6=12,18,24,30,36,42,48

OpenStudy (swissgirl):

9=9,18,27,36,45,54 6=6,12,18,24,30,36,42,48 These r the multiples

OpenStudy (anonymous):

can we start from this point?

OpenStudy (swissgirl):

yes so what is the LCM?

OpenStudy (anonymous):

Is it 18 or 36

OpenStudy (swissgirl):

Well what is the lowest number?

OpenStudy (anonymous):

18

OpenStudy (swissgirl):

Correct

OpenStudy (anonymous):

The lowest number is 18

OpenStudy (anonymous):

Now this is where I'm still getting confused.. My next next.. what should it be? 4/9 +?

OpenStudy (anonymous):

I'm not sure what to do with the 9 and the 6. Should the nenew fraction be 4/9 +9/9

OpenStudy (swissgirl):

okkk i am back lets continue

OpenStudy (swissgirl):

WE need to get our denominator to 18 right so 9*_=18 9 times what equals 18?

OpenStudy (swissgirl):

9*2=18 correct?

OpenStudy (swissgirl):

\(\huge \frac{4} {9} *\frac{2} {2}+ \frac{5} {6}* \frac{3} {3} \) \( \huge \frac{8} {18} + \frac{15} {18}=\frac{23} {18} \)

OpenStudy (anonymous):

I am so confused.. you and the other person was talking to me at the same tme.. now I have two different answere.. yours is 15/18 and the oter one is 23/18

OpenStudy (swissgirl):

mine is 23/18 too

OpenStudy (swissgirl):

look

OpenStudy (anonymous):

Oh I see it now.. but I have no idea how you got that answer.. I understood you all the way to finding the LCM.. then you left, and I couldn't find the notes. Was the LCM 18

OpenStudy (swissgirl):

ok let me get u the notes

OpenStudy (swissgirl):

http://openstudy.com/users/swissgirl#/updates/5060bcf6e4b02e139410fac4 here is the link to ur previous question

OpenStudy (anonymous):

Thanksks, I finally got to see all the notes.

OpenStudy (anonymous):

Thanks for the notes, I went back and re-read them,and I understand a little more.. however I am still a little confused as to where you got the 3 from. I'm thinking its the multiple of 3x9...and that's why you used it for the other fraction. This is the part, that I'm so confused about. not sure which numbers to use to make the other fractions with, when the denominators aren't the same.